OpenStudy (anonymous):

How do I convert the equation of this circle from general form to standard form? 2x^2 +2y^2 +4x -8y -22 =0 (I know I must complete the square, but I am stuck...)

OpenStudy (anonymous):

standard form is (x-h)^2 + (y-k)^2= r^2 I know this... how should I start?

OpenStudy (anonymous):

2x^2 +4x +___ +2y^2 -8y + ___ -22 =0 ???

OpenStudy (anonymous):

can someone help please?

OpenStudy (jdoe0001):

\(\bf 2x^2 +2y^2 +4x -8y -22 =0\\ \quad \\p \textit{let us first group, yes}\\ \quad \\ (2x^2 +4x)+(2y^2 -8y) -22 =0\\ \textit{let's take common factors}\\ 2(x^2 +2x)+2(y^2 -4y) -22 =0\\ \quad \\ \textit{now let's complete the square}\\ 2(x^2 +2x+\square^2)+2(y^2 -4y+\square^2) -22 =0\)

OpenStudy (jdoe0001):

so... what do you think our values would be?

OpenStudy (anonymous):

hello, thanks for the help! i'm reading this as we speak, so bear with me!

OpenStudy (anonymous):

completing the square means taking the first coefficient of the x^2 and y^2 values, right? so we would take 1 and divide it by 2 and then square it? so the value for each would be 1/4?

OpenStudy (jdoe0001):

darn... typos.. anyhow

OpenStudy (jdoe0001):

so, \(\bf \textit{now let's complete the square}\\ 2(x^2 +2x+\square^2)+2(y^2 -4y+\square^2) -22 =0\\ \quad \\ 2(x^2 +2(x)(\square)+\square^2)+2(y^2 -2(y)(2)+\square^2) -22 =0\) so... what do you think are our missing values?

OpenStudy (jdoe0001):

completing the square means... making it like a perfect square trinomial

OpenStudy (jdoe0001):

the middle term in a "perfect square trinomial" is 2 times the non-squared version of the other terms

OpenStudy (anonymous):

ok, but I am still confused I think. I'll look at it a bit more and let you know if I've got anything.

OpenStudy (jdoe0001):

\(\huge a^2+2ab+b^2\implies (a+b)^2\)

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=xGOQYTo9AKY

OpenStudy (anonymous):

so... the first would be: 2(x^2 + 2x +4)? the second: 2( y^2 -4y + 16) ?

OpenStudy (anonymous):

That wasn't correct then? I'll browse this video if you could keep thinking up a way to help explain it to me.